OpenStudy (anonymous):

If Log 4 (x) = 12, then log 2 (x / 4) is equal to

OpenStudy (anonymous):

@AlexandervonHumboldt2 @Nnesha

OpenStudy (anonymous):

I think it is 22 but I'm not sure

OpenStudy (anonymous):

@AlexandervonHumboldt2 is it 22

OpenStudy (michele_laino):

hint: if \[\large Lo{g_4}x = 12\] then: what is x?

OpenStudy (anonymous):

3

OpenStudy (michele_laino):

we have to apply the definition of logarithm

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

so what is the def.

OpenStudy (michele_laino):

so, we have: that definition is: the logarithm of a number is the exponent to which we have to poer the base of that logarithm in order to get that number

OpenStudy (michele_laino):

oops..power the base

OpenStudy (anonymous):

so 3

OpenStudy (michele_laino):

no, example: if we have: \[\large Lo{g_4}N = A\] then applying the definition of logarithm, we get: \[\large N = {4^A}\]

OpenStudy (anonymous):

okay

OpenStudy (michele_laino):

now, we have: \[\large Lo{g_4}x = 12\] so, what is x?

OpenStudy (michele_laino):

we have to apply the formula above

OpenStudy (anonymous):

okay so it is 22

OpenStudy (michele_laino):

explanation: I apply the definition of logarithm and I get: \[\large x = {4^{12}}\]

OpenStudy (michele_laino):

please tell me when I may continue

OpenStudy (anonymous):

48

OpenStudy (michele_laino):

no, since we can write: \[\large \frac{x}{4} = \frac{{{4^{12}}}}{4} = ...?\]

OpenStudy (michele_laino):

what is: \[\large \frac{{{4^{12}}}}{4} = ...?\]

OpenStudy (anonymous):

ummm...

Nnesha (nnesha):

use ur calculator

OpenStudy (michele_laino):

hint: \[\large \frac{x}{4} = \frac{{{4^{12}}}}{4} = {4^{11}} = {\left( {{2^2}} \right)^{11}} = {2^{22}}\]